OpenStudy (anonymous):

What is the molarity of the H+ ion when 11.1g HNO3 are dissolved in 900 ml of H2O? I'm not sure how to get the answer. Could someone please show me how to do this problem step-by-step?

OpenStudy (anonymous):

first you must find mole of HNO3 mole = mass/Mr = 11.1/63 = 0.176 volume of water 900 ml = 0.9 litres Molar = mole/volume = 0.176/0.9 = 0.196

OpenStudy (anonymous):

then you ionized HNO3 --> H+ + NO(3-) molarity of H+ equal to molarity of HNO3

OpenStudy (anonymous):

sorry, where did you get the 63 from? Mr? Did you mean M?

OpenStudy (anonymous):

Mr is realtive molecular mass H = 1 N = 14 O = 16 Mr for HNO3 = 1x1 + 1x14 + 3x16 = 63

OpenStudy (anonymous):

Oh okay. Just to clarify, after you calculate everything else and you do mol/volume then that's my final answer right?

OpenStudy (anonymous):

it depends. for this problem because only 1 ion of H+ so thats the final answer. But if there is more ion H+ you must multiply with the number of H+

OpenStudy (anonymous):

Okay, so if it had been H2NO3 then I would multiply by 2. Got it. Thank you very much!

OpenStudy (anonymous):

yup you r right. Ur welcome